# Long-term prediction of El Niño-Southern Oscillation using reservoir computing with data-driven realtime filter

### v1

@@ Abstract
Lead paragraph: Predicting long-term climate phenomena, such as the El Niño–Southern Oscillation (ENSO), is crucial for mitigating the impacts of extreme weather events worldwide on agriculture and economics. Yet, forecasting the complex systems remains a challenge, especially when different time scales of variability overlap. Furthermore, in past studies which use convectional smoothing technique, information from the future is introduced to some extent. In this paper, we introduce a new type of filter that is applicable to realtime operational forecast and combine it with a machine-learning framework known as reservoir computing, a brain-inspired method for dynamical systems modeling. An important strength of our method is that it relies solely on past data, enabling realtime predictions without the need for future observations. The parameters of the realtime filter as well as the hyperparameters of the resevoir computing model are systematically calibrated with the Bayesian optimizations. Using this method, we show that the prediction horizon of ENSO can be extended up to two years, which is longer than most of prediction horizons previously reported. While we focus on ENSO forecasts in this article, our method can also be applied to a wide range of complex, high-dimensional phenomena that demand robust and efficient long-term forecasting.
@@ I Introduction
The reservoir computing is a brain-inspired machine-learning technique that employs a data-driven dynamical system Jaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). It is a method effective in predicting time series and frequency spectra in chaotic behaviors and is capable of learning time series of high-dimensional dynamical systems such as fluid flow  Nakai and Saiki (2018, 2021); Kobayashi et al. (2021); Nakai and Saiki (2024). In particular, Nakai and Saiki Nakai and Saiki (2021) clarified that delay-coordinates with appropriate delay time and dimension of the delay-coordinates are efficient when the number of observable variables is smaller than the effective dimension of the attractor. The reservoir computing can predict the statistical quantity of a particular macroscopic variable of chaotic fluid flow, which cannot be calculated from a direct numerical simulation of the Navier–Stokes equation because of its high computational cost Kobayashi et al. (2021). The reservoir computing approach has been proven to be effective in prediction of complex atmospheric systems such as Madden-Julian Oscillation Suematsu et al. (2022), the Asian summer monsoon Mitsui and Boers (2021), North Atlantic Oscillation Huang and Fu (2023) as well as ENSO Hassanibesheli, Kurths, and Boers (2022); Guardamagna et al. (2024).
@@
The aim of this study is to establish a general methodology for performing long-term prediction of various time series in operational realtime formulation, which excludes information from the future. Applying the band-pass filter to the time series data is proven to be effective for extending the prediction horizon of the ENSO time series. For a robust long-term prediction, we introduce a new data-driven filter. It is known that the high-frequency disturbances degrade the performance of the machine learning prediction, while the low-frequency behavior such as trends and multidecadal variability make construction of machine-learning model difficult Hassanibesheli, Kurths, and Boers (2022). In past studies, the formulation of predicting ENSO by artificial neural networks requires the application of some convectional bandpass filter techniques such as moving averageGuardamagna et al. (2024) or Butterworth filter Hassanibesheli, Kurths, and Boers (2022); Ham, Kim, and Luo (2019), which incorporate information from the future time steps into the filtered time series. Strictly speaking, even the convectional Niño indices include future information by the moving average. Therefore, those methodologies are not practical to be applied in a realtime operational forecast. This study is the first successful attempt to develop a filtering method that extracts signals with cycles of 3 to 8 years using only past time series. For the filtered Niño-3.4 time series data, we construct a data-driven model. In this study, we employ reservoir computing to construct a machine-learning model capable of predicting the multi-year dynamics of the ENSO for over two years.
@@ II Data and Methods
II.1 Realtime fitering method
In this subsection, we present a new method to find a suitable filter for the corresponding time series data. Note that the corresponding Python codes of realtime filter and reservoir computing combining with Bayesian optimization are available in the repository Yoneda (2024). The pure reservoir computing code is also available.Tanaka, Nakane, and Hirose (2021) To find the appropriate filter, we use Bayesian optimization and the following weight function of the moving average:
@@ II.1 Realtime fitering method
where $d_{1},d_{2},r_{1},r_{2},c,w>0$ are the parameters of the corresponding objective function (3). Let $\Omega\subset\mathbb{Z}$ be a prescribed finite sequence. Having obtained $\Psi$ by the optimization, then we applied the weighted moving average to the original data $y:\Omega\to\mathbb{R}$ as follows:
@@
We emphasize that since the weight function $\Psi$ is localized in the timeline, in particular, it does not contain any future information, this bandwidth cannot have a unique threshold due to the Fourier uncertainty principle. Here is a rough explanation of the bandwidth. If $c$ is close to zero, then the Fourier transform of the corresponding weight function is close to the sinc function, which represents narrow bandwidth (compared to the case when $c$ is large). To the contrary, if $c$ tends to infinity, then the Fourier transform of the corresponding weight function tends to identically one (since $\Psi$ tends to the Dirac-delta function), which represents broad bandwidth (compared to the case when $c$ is small). The main philosophy of this filter method is to effectively reduce the “indefinite factor” (such as white noise) without using the conventional Fourier transform, since the time domain of the training data $\Omega$ is, of course, always finite, and, applying the Fourier transform to such finite elements is no longer mathematically rigorous. To effectively reduce the indefinite factor, we use the idea of “dictionary”, in other word, “key-value pairs”. First, we discretize the range of the filtered data $y^{*}$ as follows: For an integer $K$ which is greater than $1$, we choose $\{a_{k}\}_{k=1}^{K}\subset\mathbb{R}$ such that $a_{1}<a_{2}<\cdots<a_{K}$.
@@
and we maximize the following objective function under the condition that the autocorrelation between the original and the filterd data is not small:
@@
where $\gamma\in(0,1]$ is the match rate of $a_{k}$ for each $\sigma_{n}^{L}$. Here, $\gamma$, as well as $L_{max}$ and $L_{min}$ are prescribed parameters in the numerical implementation. $\gamma$ is close to 1. $L_{max}$ and $L_{min}$ are determined so that the length of frequent patterns are sufficiently covered.
II.2 Data description
We use monthly sea surface temperature (SST) for the 1870–2022 period, retrieved from the Hadley Centre Global Sea Ice and Sea Surface Temperature dataset Centre (2006). The one-degree latitude-longitude grid data is averaged over the $5^{\circ}$–$5^{\circ}$, $170^{\circ}$–$5^{\circ}$ regions in the tropical central and eastern Pacific, which are commonly used to define the Niño-3.4 index. To obtain the SST anomaly from the typical seasonal cycle, the 1971–2000 monthly climatological values are subtracted. We then filter the unrequired frequency ranges by the realtime filter described in the previous subsection. The convectional Niño-3.4 index uses 5-month running mean to define El Niño and La Niña events. However, in this study, we utilize monthly SST values without applying any moving average to exclude future information and enable realtime prediction. Hereafter, we refer to this filtered time series as the realtime SST anomaly.
@@ II.2 Data description
II.3 Reservoir computing
The echo state network is a widely used reservoir computing architecture based on a recurrent neural network, consisting of an input layer, a reservoir state vector, and an output layerJaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). A vector of an observed variable as input and a reservoir state vector at time step $t$ is updated as
where $\mathbf{u}(t)\in\mathbb{R}^{M}$ is the input vector of the time series data, $\mathbf{r}(t)\in\mathbb{R}^{N}~{}(N\gg M)$ is the reservoir state vector, $\Delta t$ represents the time interval at which the time series is updated (1 month in this study), and $\sigma_{\text{in}}$ is the scaling parameter of the input matrix $W_{\text{in}}$. $\mathbf{W}_{\text{in}}\in\mathbb{R}^{N\times M}$ and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{M\times N}$ are input and output matrices; $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a matrix that depicts the recurrent connectivity within the reservoir. The density of the non-zero elements of $\mathbf{A}$ is determined by the parameter $p$. The maximum eigenvalue of $\mathbf{A}$ is controlled by the parameter $\rho$ $(|\rho|<1)$ to satisfy the echo state property that affects the performance of the data-driven model. $\alpha$ ($0<\alpha\leq 1$) is the coefficient called leaking rate that adjusts the smoothness of the time evolution of $\mathbf{r}(t)$. Hyperbolic tangent function for a vector $\mathbf{q}=(q_{1},q_{2},\ldots,q_{N})^{\text{T}}$ is defined as $\tanh(\mathbf{q})=(\tanh(q_{1}),\tanh(q_{2}),\ldots,\tanh(q_{N}))^{\text{T}}$ where superscript T represents the transpose.
@@ II.3 Reservoir computing
Let $y^{*}$ be the realtime SST anomaly. For these reasons above, we introduce an $M$-dimensional delay-coordinate vector of realtime SST anomaly with a delay-time $\Delta\tau$ as input and output data $\mathbf{u}$ in Eq. (4):
@@ II.4 Bayesian optimization of hyperparameters
For both the realtime filter and the reservoir model, the combinations of hyperparameters are selected using Bayesian optimization with Optuna Akiba et al. (2019). The hyperparameters of the realtime filter are chosen to maximize the product of the objective function in Eq. (3) when applied to the monthly SST anomaly, and the maximum lag-correlation between the original and filtered monthly SST anomaly. The hyperparameters of the reservoir computing model, namely, $M$, $\Delta\tau$, $N$, $\beta$, $p$, $\rho$, $\alpha$, and $\sigma_{\text{in}}$, as well as the random matrices $\mathbf{A}$ and $\mathbf{W_{\text{in}}}$, are chosen based on the prediction performance of 120 sequences of training and predictions whose $t_{0}$ ranges from January 1986 to December 1995. As a measure of the model performance, we use the all-season correlation skill $C(\mu)$
$C(\mu)=\sum_{m=1}^{12}\frac{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})(P_{\eta,m,\mu}-\overline{P_{m,\mu}})}{\sqrt{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})^{2}\sum_{\eta=\eta_{s}}^{\eta_{e}}(P_{\eta,m,\mu}-\overline{P_{m,\mu}})^{2}}}$ (7)
as a function of the forecast lead months $\mu$. Here, $Y$ and $P$ denote the observed and the predicted values, respectively. $\overline{Y_{m}}$ and $\overline{P_{m,l}}$ denote the climatologies with respect to the calendar month $m$ (from 1 to 12) and the forecast lead months $\mu$. The label $\eta$ denotes the forecast target year. $\eta_{s}$ and $\eta_{e}$ denote the earliest (that is, 1986) and the latest year (that is, 1995) of the period, respectively. $C(\mu=24)$ is maximized here. The determination of the random matrices $\mathbf{W}_{\text{in}}$ and $\mathbf{A}$ is as follows. First, using two consecutive integers $i$ and $i+1$ as seed values, two $N\times N$ matrices $\mathbf{W}_{i}$ and $\mathbf{W}_{i+1}$ are generated with a density of non-zero elements equal to $p$. Next, by linearly interpolating between the two matrices while maintaining the density $p$, a matrix $\mathbf{W}_{i+x}$ characterized by a real seed value $x$ $(0<x<1)$ is created. Finally, $\mathbf{W}_{i+x}$ is scaled to match the condition of spectral radius $\rho$.
@@ III.1 Performance of the new filter
By applying the new realtime filter, we obtain a target time series that preserves information on long-term ENSO variations while enhancing the predictive performance of the reservoir computing model. The filtered realtime SST anomaly is shown in Fig.4a along with the original time series. Due to the asymmetry of the filter function with respect to the positive and negative time ranges, the overall time series shifts backward by approximately five months. The shift is highlighted by the lag correlation between the original and the filtered time series (Fig.4b). The maximum correlation is 0.837 with a lag of five months, and we assume that the filtered realtime SST anomaly is can be used as an alternative index for the state of ENSO.
@@ III.2 Prediction skill of the model
As a result of the Bayesian optimization, the optimal set of the hyper parameters for reservoir computing model is obtained as listed in Table 2. To evaluate the prediction skill of the current method, the sequence of training and prediction is conducted 180 times with different $t_{0}$ every month between January 2001 and December 2015. The forecast skill of the filtered realtime SST anomaly was evaluated using the all-season correlation skill in Eq. (7) with $\eta_{s}=2001$ and $\eta_{e}=2015$.
@@
We focus on enhancing our data-driven model’s ability to predict the long-term behavior of ENSO over timescales beyond one year. With the optimal hyper-parameter settings of the realtime filter and the reservoir computing model, the average all-season correlation skill remains above 0.5 for 29 months (Fig. 5). Even discounting the fact that the filtered time series to be modeled is shifted about 5 months into the future due to the asymmetry of the filter function, the performance of the current model succeeds in predicting the dynamics of ENSO for 24 months (i.e., two years). The two-year prediction horizon achieved in this study surpasses the 21 months reported in Hassanibesheli et al. (2022)Hassanibesheli, Kurths, and Boers (2022) using reservoir computing, despite we strictly avoid the use of future information. Furthermore, our result is comparable to the two-year prediction horizon obtained with a deep neural network model Wang et al. (2024), while it should be noted that the target of prediction in this study is the filtered time series.
@@
Figure 5: Prediction skill of the data driven model. All-season correlation skill for ecch forecast lead time is obtained by averaging over 180 predictions between January 2001 and December 2015 based on Eq. (7). The horizontal red dashed line denotes a correlation of 0.5, displayed as a benchmark for the model’s predictive capability.
@@
Figure 6: Comparison between the predicted and the target time series of filtered realtime SST anomaly. (a) 180 predictions and (b) corresponding target. Each row in the two-dimensional plot corresponds to an individual forecast. The horizontal axis indicates the lead time from the start of the prediction, while the vertical axis shows the prediction start time step, with January 2001 set as zero. The maximum forecast lead time is 36 months.
The predicted time series are displayed along with the target in Fig. 6. In several successful cases (e.g. start month 85 and 110), we have been able to capture the major positive and negative peaks of the realtime SST anomaly up to three years in advance.

### v2

@@ Abstract
Lead paragraph: Predicting long-term climate phenomena, such as the El Niño–Southern Oscillation (ENSO), is crucial for mitigating the impacts of extreme weather events worldwide on agriculture and economics. Yet, forecasting the complex systems remains a challenge, especially when different time scales of variability overlap. Furthermore, in past studies which use conventional smoothing technique, information from the future is introduced to some extent. In this paper, we introduce a new type of filter that is applicable to realtime operational forecast and combine it with a machine-learning framework known as reservoir computing, a brain-inspired method for dynamical systems modeling. An important strength of our method is that it relies solely on past data, enabling realtime predictions without the need for future observations. The parameters of the realtime filter as well as the hyperparameters of the reservoir computing model are systematically calibrated with the Bayesian optimizations. Using this method, we show that the prediction horizon of ENSO can be extended up to two years. While we focus on ENSO forecasts in this article, our method can also be applied to a wide range of complex, high-dimensional phenomena that demand robust and efficient long-term forecasting.
@@ I Introduction
Reservoir computing is a brain-inspired machine-learning technique that employs a data-driven dynamical system Jaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). It is a method effective in predicting time series and frequency spectra in chaotic behaviors and is capable of learning time series of high-dimensional dynamical systems such as fluid flow  Nakai and Saiki (2018, 2021); Kobayashi et al. (2021); Nakai and Saiki (2024). In particular, Nakai and Saiki Nakai and Saiki (2021) clarified that delay-coordinates with appropriate delay time and dimension of the delay-coordinates are efficient when the number of observable variables is smaller than the effective dimension of the attractor. The reservoir computing can predict the statistical quantity of a particular macroscopic variable of chaotic fluid flow, which cannot be calculated from a direct numerical simulation of the Navier–Stokes equation because of its high computational cost Kobayashi et al. (2021). The reservoir computing approach has been found to be effective in prediction of complex atmospheric systems such as Madden-Julian Oscillation Suematsu et al. (2022), the Asian summer monsoon Mitsui and Boers (2021), North Atlantic Oscillation Huang and Fu (2023) as well as ENSO Hassanibesheli, Kurths, and Boers (2022); Guardamagna et al. (2024).
@@
The aim of this study is to establish a general methodology for performing long-term prediction of various time series in operational realtime formulation, which excludes information from the future. In this study, applying the band-pass filter to the time series data is demonstrated to be effective for extending the prediction horizon of the ENSO time series. For a robust long-term prediction, we introduce a new data-driven filter. It is known that high-frequency disturbances degrade the performance of machine learning prediction, while low-frequency behavior, such as trends and multidecadal variability, makes the construction of machine-learning models difficult Hassanibesheli, Kurths, and Boers (2022). In past studies, the formulation of predicting ENSO by artificial neural networks requires the application of some conventional bandpass filter techniques such as moving averageGuardamagna et al. (2024) or Butterworth filter Hassanibesheli, Kurths, and Boers (2022); Ham, Kim, and Luo (2019), which incorporate information from the future time steps into the filtered time series. Strictly speaking, even the conventional Niño indices include future information by the moving average. Therefore, those methodologies are not practical to be applied in a realtime operational forecast. This study is the first successful attempt to develop a filtering method that extracts signals with cycles of 3 to 8 years using only past time series. For the filtered Niño-3.4 time series data, we construct a data-driven model. In this study, we employ reservoir computing to construct a machine-learning model capable of predicting the multi-year dynamics of the ENSO for over two years.
@@ II Data and Methods
II.1 Data description
We use monthly sea surface temperature (SST) for the 1870–2022 period, retrieved from the Hadley Centre Global Sea Ice and Sea Surface Temperature dataset Centre (2006). The one-degree latitude-longitude grid data is averaged over the $5^{\circ}$–$5^{\circ}$, $170^{\circ}$–$120^{\circ}$ regions in the tropical central and eastern Pacific, which are commonly used to define the Niño-3.4 index. To obtain the SST anomaly from the typical seasonal cycle, the 1971–2000 monthly climatological values are subtracted. Hereafter, we refer to this time series as the realtime SST anomaly. The conventional Niño-3.4 index uses 5-month running mean to define El Niño and La Niña events. However, in this study, we utilize monthly SST values without applying any moving average to exclude future information and enable realtime prediction. We then filter the unrequired frequency ranges by the realtime filter described in the next subsection.
II.2 Realtime fitering method
In this subsection, we present a new method to find a suitable filter for the corresponding time series data. Note that the corresponding Python codes of realtime filter and reservoir computing combined with Bayesian optimization are available in the repository Yoneda (2024). The pure reservoir computing code is also available.Tanaka, Nakane, and Hirose (2021) To find the appropriate filter, we use Bayesian optimization and the following weight function of the moving average:
@@ II.2 Realtime fitering method
where $d_{1},d_{2},r_{1},r_{2},c,w>0$ are the parameters of the corresponding objective function (3). In what follows, we explain how to obtain this objective function. Let $\Omega\subset\mathbb{Z}$ be a prescribed finite sequence. Having obtained $\Psi$ by the optimization, we then applied the weighted moving average to the original data $y:\Omega\to\mathbb{R}$ as follows:
@@
We emphasize that since the weight function $\Psi$ is localized in the timeline, in particular, it does not contain any future information, this bandwidth cannot have a unique threshold due to the Fourier uncertainty principle. Here is a rough explanation of the bandwidth. If $c$ is close to zero, then the Fourier transform of the corresponding weight function is close to the sinc function, which represents a narrow bandwidth (compared to the case when $c$ is large). To the contrary, if $c$ tends to infinity, then the Fourier transform of the corresponding weight function tends to identically one (since $\Psi$ tends to the Dirac-delta function), which represents broad bandwidth (compared to the case when $c$ is small). The main philosophy of this filter method is to effectively reduce the “indefinite factor” (such as white noise) without using the conventional Fourier transform, since the time domain of the training data $\Omega$ is, of course, always finite, and, applying the Fourier transform to such finite elements is no longer mathematically rigorous. To effectively reduce the indefinite factor, we use the idea of “dictionary”, in other word, “key-value pairs”. First, we discretize the range of the filtered data $y^{*}$ as follows: For an integer $K$ which is greater than $1$, we choose $\{a_{k}\}_{k=1}^{K}\subset\mathbb{R}$ such that $a_{1}<a_{2}<\cdots<a_{K}$.
@@
and we maximize the following objective function under the condition that the autocorrelation between the original and the filtered data is not small:
@@
where $\gamma\in(0,1]$ is the match rate of $a_{k}$ for each $\sigma_{n}^{L}$. Here, $\gamma$, as well as $L_{max}$ and $L_{min}$ are prescribed parameters in the numerical implementation. $\gamma$ is close to 1. $L_{max}$ and $L_{min}$ are determined so that the length of frequent patterns is sufficiently covered.
II.3 Reservoir computing
The echo-state network (ESN) is a widely used reservoir computing architecture based on a recurrent neural network, consisting of an input layer, a reservoir state vector, and an output layerJaeger (2001); Jaeger and Haas (2004); Lu et al. (2017); Lu, Hunt, and Ott (2018). A vector of an observed variable as input and a reservoir state vector at time step $t$ is updated as
where $\mathbf{u}(t)\in\mathbb{R}^{M}$ is the input vector of the time series data, $\mathbf{r}(t)\in\mathbb{R}^{N}~{}(N\gg M)$ is the reservoir state vector, $\Delta t$ represents the time interval at which the time series is updated (1 month in this study), and $\sigma_{\text{in}}$ is the scaling parameter of the input matrix $\mathbf{W_{\text{in}}}$. $\mathbf{W}_{\text{in}}\in\mathbb{R}^{N\times M}$ and $\mathbf{W}_{\text{out}}\in\mathbb{R}^{M\times N}$ are input and output matrices; $\mathbf{A}\in\mathbb{R}^{N\times N}$ is a matrix that depicts the recurrent connectivity within the reservoir. The density of the non-zero elements of $\mathbf{A}$ is determined by the parameter $p$. The maximum absolute value of the eigenvalue of $\mathbf{A}$ is called the spectral radius and denoted by $\rho$. The spectral radius is related to the echo state property (ESP), which means that the effect of initial conditions vanishes after a certain amount of time. Setting the spectral radius to a value less than unity has been used as a rule of thumb to ensure that the model possesses the ESP. However, ESP often holds for $\rho>1$ Lukoševičius (2012). Yildiz et al. (2012) demonstrated that no universally applicable rule exists between the ESP and the optimal spectral radiusYildiz, Jaeger, and Kiebel (2012). It should be determined through task-specific experimentation, taking into account the nature of the driving input. In the present study, we impose the constraint $\rho<1$ with the intention of limiting the parameter search space in Bayesian optimization, thereby facilitating the identification of the optimal value. $\alpha$ ($0<\alpha\leq 1$) is the coefficient called leaking rate that adjusts the smoothness of the time evolution of $\mathbf{r}(t)$. Hyperbolic tangent function for a vector $\mathbf{q}=(q_{1},q_{2},\ldots,q_{N})^{\text{T}}$ is defined as $\tanh(\mathbf{q})=(\tanh(q_{1}),\tanh(q_{2}),\ldots,\tanh(q_{N}))^{\text{T}}$ where superscript T represents the transpose.
@@ II.3 Reservoir computing
Let $y^{*}$ be the filtered realtime SST anomaly. For these reasons above, we introduce an $M$-dimensional delay-coordinate vector of filtered realtime SST anomaly with a delay-time $\Delta\tau$ as input and output data $\mathbf{u}$ in Eq. (4):
@@ II.4 Bayesian optimization of hyperparameters
For both the realtime filter and the reservoir model, the combinations of hyperparameters are selected using Bayesian optimization with Optuna Akiba et al. (2019). The hyperparameters of the realtime filter are chosen to maximize the product of the objective function in Eq. (3) when applied to the realtime SST anomaly, and the maximum lag-correlation between the original and filtered monthly SST anomaly. The hyperparameters of the reservoir computing model, namely, $M$, $\Delta\tau$, $N$, $\beta$, $p$, $\rho$, $\alpha$, and $\sigma_{\text{in}}$, as well as the random seed to generate the matrices $\mathbf{A}$ and $\mathbf{W_{\text{in}}}$, are chosen based on the prediction performance of 120 sequences of training and predictions whose $t_{0}$ ranges from January 1986 to December 1995. As a measure of the model performance, we use the all-season correlation skill $C(\mu)$
$C(\mu)=\frac{1}{12}\sum_{m=1}^{12}\frac{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})(P_{\eta,m,\mu}-\overline{P_{m,\mu}})}{\sqrt{\sum_{\eta=\eta_{s}}^{\eta_{e}}(Y_{\eta,m}-\overline{Y_{m}})^{2}\sum_{\eta=\eta_{s}}^{\eta_{e}}(P_{\eta,m,\mu}-\overline{P_{m,\mu}})^{2}}}$ (7)
as a function of the forecast lead months $\mu$. Here, $Y$ and $P$ denote the observed and the predicted values, respectively. $\overline{Y_{m}}$ and $\overline{P_{m,l}}$ denote the climatologies with respect to the calendar month $m$ (from 1 to 12) and the forecast lead months $\mu$. The label $\eta$ denotes the forecast target year. $\eta_{s}$ and $\eta_{e}$ denote the earliest (that is, 1986) and the latest year (that is, 1995) of the period, respectively. $C(\mu=24)$ is maximized here. The determination of the random matrix $\mathbf{A}$ is as follows. First, using two consecutive integers $i$ and $i+1$ as seed values, two $N\times N$ matrices $\mathbf{A}_{i}$ and $\mathbf{A}_{i+1}$ are generated with a density of non-zero elements equal to $p$. Next, by linearly interpolating between the two matrices while maintaining the sparsity $p$, a matrix $\mathbf{A}_{i+x}$ characterized by a real seed value $x$ $(0<x<1)$ is created. Finally, $\mathbf{A}_{i+x}$ is rescaled to match the condition of spectral radius $\rho$. For the determination of the random matrix $\mathbf{W}_{\text{in}}$, it is similarly defined as a matrix obtained by linearly interpolating between two random matrices $\mathbf{W}_{i}$ and $\mathbf{W}_{i+1}$.
@@ III.1 Performance of the new filter
By applying the new realtime filter, we obtain a target time series that preserves information on long-term ENSO variations while enhancing the predictive performance of the reservoir computing model. The filtered realtime SST anomaly is shown in Fig.4a along with the original time series. Due to the asymmetry of the filter function with respect to the positive and negative time ranges, the overall time series shifts backward by approximately five months. The shift is highlighted by the lag correlation between the original and the filtered time series (Fig.4b). The maximum correlation is 0.837 with a lag of five months, and we assume that the filtered realtime SST anomaly can be used as an alternative index for the state of ENSO.
@@ III.2 Prediction skill of the model
As a result of Bayesian optimization, the optimal set of hyperparameters for the reservoir computing model is obtained as listed in Table 2. To evaluate the prediction skill of the current method, the sequence of training and prediction is conducted 180 times with different $t_{0}$ every month between January 2001 and December 2015. The prediction skill of the filtered realtime SST anomaly was evaluated using the all-season correlation skill in Eq. (7) with $\eta_{s}=2001$ and $\eta_{e}=2015$.
@@
We focus on enhancing our data-driven model’s ability to predict the long-term behavior of ENSO over timescales beyond one year. The definition of successful prediction is based on maintaining a correlation skill above 0.5 for a given period, which is a criterion commonly used in previous studiesHassanibesheli, Kurths, and Boers (2022); Ham, Kim, and Luo (2019); Wang et al. (2024). With the optimal hyper-parameter settings of the realtime filter and the reservoir computing model, the average all-season correlation skill remains above 0.5 for 29 months (red line in Fig. 5). Even discounting the fact that the filtered time series to be modeled is shifted about 5 months into the future due to the asymmetry of the filter function, the performance of the current model succeeds in predicting the dynamics of ENSO for 24 months (i.e., two years). Note that in this case, the predictions are made using a fixed set of $\mathbf{A}$ and $\mathbf{W_{\text{in}}}$ that maximizes the objective function during the model selection process. In Fig. 5, the correlation skill of the same evaluation period is also displayed for the cases where 100 sets of $\mathbf{A}$ and $\mathbf{W_{\text{in}}}$ are randomly selected (blue solid line for the mean and shading for the standard deviation). From this figure, it can be observed that the optimization in the model selection process allows for extracting optimal random matrices that result in a reservoir model with outstanding prediction skill in the model evaluation process over a different data period.
@@
Figure 5: Prediction skill of the data-driven model. All-season correlation skill for ecch forecast lead time is obtained from 180 predictions between January 2001 and December 2015 based on Eq. (7). The red line shows the prediction obtained using the optimal matrices $\mathbf{W}_{\text{in}}$ and $\mathbf{A}$, which were determined simultaneously with the hyperparameters during the model selection process. The blue solid line represents the average of the correlation skill when randomly generated 100 combination of $\mathbf{W}_{\text{in}}$ and $\mathbf{A}$ are used with the same hyperparameters as in Table 2, while the blue shading indicates the range of ±1 standard deviation. The horizontal gray dashed line denotes a correlation of 0.5, displayed as a benchmark for the model’s prediction capability.
The main achievement of this study is the machine learning prediction of the long-term dynamics of ENSO, with a lead time up to two years, strictly avoiding the use of future information by the realtime filter. The two-year prediction horizon is achieved using the simpler architechture of reservoir computing, compared to the deep neural network model by Wang et al. (2024)Wang et al. (2024), while it should be noted that the target of prediction in this study is the filtered time series.
@@
Figure 6: Comparison between the predicted and the target time series of filtered realtime SST anomaly. (a) The reference target and (b) the 180 runs of prediction. Each column in the two-dimensional plot corresponds to an individual run of 36-month prediction. The vertical axis indicates the lead time from the start of the prediction, while the horizontal axis shows the prediction start time step, with January 2001 set as zero. (c) Comparison of the prediction versus the target filtered realtime SST anomaly for different lead months: 6 months (blue line), 12 months (green line), and 24 months (orange line).
The predicted time series (Fig. 6b) are displayed along with the corresponding target (Fig. 6a). In some successful cases (e.g. those that start during 2010 and 2011), we have been able to capture the major positive and negative peaks of the filtered realtime SST anomaly up to about three years in advance. The predictions for different lead times (Fig. 6c) show a high visual similarity to the target (filtered realtime SST anomaly). In particular, even the prediction with a lead time of 24 months faithfully captures the large-amplitude cycle of ENSO observed from 2010 to 2014. It should also be noted that the long-term predictions of the model in this study show relatively low accuracy in representing the signal of the record-breaking strong El Niño that occurred between 2015 and 2016.